# Towards Identifiability of Interventional Stochastic Differential Equations

### v2

@@ 5 Experiments
To validate the theory given above, we consider experiments demonstrating the recovery of SDE parameters and generalization to unseen interventions. To show the broad applicability of the theory, we consider multiple loss functions: a loss acting directly on the parameters in the linear case [33], a kernelized Stein discrepancy [3], and rollout loss for neural SDEs [21].
@@ 5.1 Linear SDE Recovery
Because we are fitting linear SDEs where the stationary distribution has an essentially closed-form (up to solving the Lyapunov equation), we can choose a very simple loss to verify the claim of identifiability (Theorem 4.4). This loss is akin to the one used in Bicycle [33] that fits interventional linear SDEs. Assuming the noise scale $\epsilon$ is known and $L$ denotes the true drift, we fit the parameters $\hat{A}$, $\hat{B}$, and $\hat{D}$. Letting our estimate for the drift be $\hat{L}=\hat{A}\hat{B}-\hat{D}$, the loss function is,
$\mathcal{L}_{lin}(\hat{L})=\|\hat{L}^{-1}C-L^{-1}C\|_{F}+\|\hat{L}\omega+\omega\hat{L}^{T}+\epsilon I\|_{F}.$ (9)
In other words, the loss is an unweighted sum of the mean loss for every intervention, and a loss controlling how well the parameterized drift $\hat{L}$ induces the covariance $\omega$ that is shared among all interventions. We optimize $\hat{A}$, $\hat{B}$ and $\hat{D}$ (constrained to be diagonal) with gradient descent.
@@
We confirm our theory: when the decay is known, $r-2$ interventions fails to grant identifiability with poor performance and extremely high variance, even with many independent initializations, while $r$ interventions clearly succeed. When the decay isn’t known, we use a modest oversampling of $r\log(n)$ interventions. This amount is informed by the literature on matrix completion, where $\Omega(rn\log(n))$ revealed matrix entries were proven to be necessary for a rank $r$ matrix recovery problem [7]. We scale down by a factor of $n$ because in our setting each intervention yields the mean vector of the perturbed linear SDE in $\mathbb{R}^{n}$, hence $n$ measurements.
(a) Error in $n=100$ with fixed decay
(b) Error in $n=200$ with fixed decay
(c) Error in $n=100$ with learned decay
(d) Error in $n=200$ with learned decay
Figure 2: Normalized Frobenius error of learned drift against true drift in low-rank linear SDEs with $k$ independent Gaussian interventions. Error bars are standard deviation over 5 independent instances.
@@ 5.2 Synthetic Nonlinear SDE Generalization
To assess the insight from the theory that the parameters are identifiable with an arbitrary activation function (Theorem 4.8), we propose a simple architectural improvement wherein we learn the activation function. The theory requires the activation function be monotonic, but we consider general learnable MLPs for improved optimization. Furthermore, the simple method for extracting a GRN we apply in section 5.3 requires a two-layer network, so learning the activation is a way to increase expressiveness without sacrificing interpretability. Learnable activations [16] have been proposed before [1] but to our knowledge not previously applied to SDE parameterizations.
To assess generalizability to unseen interventions, we consider a loss function proposed by [26]: the kernel deviation from stationarity (KDS), or equivalently the kernelized Stein discrepancy of the stationary distribution [3]. When the parametric drift is defined as $v_{\theta}(x)=A\sigma(Bx)-Dx$ and the target stationary distribution is $\mu$, the KDS is,
@@
where $k$ is a given kernel, $\mathcal{A}$ is the generator the parameterized SDE defined as $\mathcal{A}h=v_{\theta}\cdot\nabla h+\frac{\epsilon}{2}\Delta h$, and $\mathcal{A}_{x}$ is the generator acting on the kernel as a function of its first element, likewise $\mathcal{A}_{x^{\prime}}$ acting on the second element. More details on the derivation of this loss are given in [26].
For evaluation, we choose another notion of distributional distance. We observed numerical instability in the Sinkhorn divergence without large entropic regularization [9]. Therefore, we follow [43, 26] and consider the mean squared error (MSE) between the true and predicted distribution. We find training on functions with monotonic activations is numerically unstable when using noise small enough to reach the limit proven in Theorem 4.8, and therefore study more complicated activations in our experiments, using sinusoidal activations. Further details are given in Section B.2.
The results are given in Table 1. We observe a clear benefit for low noise SDEs, which gets smaller as the noise gets bigger and hence further away from our proven settings. This is consistent with our theory, and also intuitive: as the noise gets larger, the stationary distrubtion gets smoother and has less dependence on the intervention vectors. Nevertheless, in the low noise regime we see that learnable activations improve prediction on test interventions without overfitting.
Table 1: Mean distance between true and predicted distribution with $n=20$, $r=3$, averaged over 20 independent seeds with 20 test interventions per seed, comparing a fixed (sigmoid) activation versus a learned activation.
@@ 5.3 Simulated GRN SDE Generalization
We consider an applied setting for learnable activations, and how they impact the recovery of GRNs. Because there are few fully known regulatory networks we rely on semi-synthetic data to produce data with similar characteristics to true perturbed transcriptomic samples and a ground-truth GRN.
We use the PerturbODE model from [25], similar to our setup and modified to handle a loss for parameterized SDEs [21]. We consider the inferred GRN as the matrix multiplication of the weight matrices in the MLP that parameterizes the drift, corresponding to a first-order Taylor approximation of the dynamics. GRN recovery can be measured as a classification task of individual edges. As a baseline, we also consider the Bicycle interventional linear SDE method [33]. To simulate a known GRN, we use the boolean ODE based-simulator BEELINE [32]. BEELINE uses the provided GRN to parametrize an SDE in the space of mRNA expression and protein expression, with physically plausible regulation dynamics. More precise details are given in the Appendix in Section B.3.
@@
Figure 3: Gene regulatory network recovery for data from BEELINE among SDE, SDE with learnable activation, and Bicycle model. Error bars are over 5 independent runs.
@@ 6 Discussion
Although we can confirm identifiability directly in the linear case because we have access to a simple loss function, direct recovery of parameters is very impractical for the nonlinear SDE parameterization. Nevertheless, the positive results with improved generalization using learned activations is good evidence to confirm the theory. It may be necessary to increase the capacity of the SDE beyond two-layer neural networks, if the extracted GRN could be proven identifiable.
We note some limitations of the the given theory. The results demonstrate identifiability in an idealized setup with infinitesimal small noise and restrictions on the possible drift functions. Generalizing to larger noise regimes may be possible with higher order moment information about the stationary distribution, although in general understanding the stationary distribution outside of linearization is a very challenging task. Considering different parameterizations of the drift or how interventions act on the SDE are especially interesting extensions for future work.
@@ 7 Conclusion
Acknowledgments
This work was made possible by support from the MacMillan Family and the MacMillan Center for the Study of the Non-Coding Cancer Genome at the New York Genome Center.
@@ Appendix A Proofs of Results
and set $D=I$ and $\epsilon=1$. It is then straightforward to confirm these choices yield a Lyapunov equation that satisfies,
@@
Furthermore, any skew-symmetric matrix $Q$ obeys the trivial Lyapunov equation $Q\omega+\omega Q^{T}=0$, so we can always add $Q\omega^{-1}$ to the vector field $L$ without changing the covariance.
Choose $Q$ to be only supported in the top left $r\times r$ block, then $AB+Q\omega^{-1}$ will still be rank $r$. And in the worst case, every intervention is of the form $e_{i}$ for $i>r$, and these can’t distinguish between different values of $Q$.
@@
Suppose the true $A$ and $B$ are sampled according to Assumption 4.2, with the constraint that a.s. $\|A\|,\|B\|\leq\sqrt{\gamma}$. Then $L=AB-D$ is a.s. identifiable.
@@
00 0 $=L\omega P+\omega L^{T}P+\epsilon P$ (21)
@@
$(\omega D+\epsilon I)P$ is rank $n-r$, so a.s. $\operatorname{im}(\omega D+\epsilon I)P\oplus\operatorname{im}C$ is $n$ dimensional. Hence we recover $L^{-1}$ and therefore $L$.
@@
Sample $A\in\mathbb{R}^{n\times r}$ and $B\in\mathbb{R}^{r\times n}$ according to Assumption 4.2. Then a.s. $AB$ is not identifiable with $r-2$ interventions.
@@
Assume the decay is fixed at $D=I$, so $L=AB-I$ with induced covariance $\omega$. With $r-2$ interventions, $ABL^{-1}C$ is at most rank $r-2$, and by assumption the kernel of $A^{T}$ is at most dimension $n-r$, so there is guaranteed to be a two dimensional subspace orthogonal to the previous spaces, say with basis vectors $u$ and $v$. Let $Q^{*}=uv^{T}-vu^{T}$.
@@
Suppose we have $n-r-1$ interventions. Since $A$ has $r$ columns, there must be a vector $u$ in $\mathbb{R}^{n}$ orthogonal to the columns of $A$ and $C$.
@@
Consider a vector field $v(x)=A\sigma(Bx)-x+c$ with $\|A\|,\|B\|\leq 1$, $\|\sigma\|_{\infty}=\gamma<1$. Let $p$ be the stationary distribution of our usual SDE, $x^{*}$ be the unique stationary point of $v$, and $P$ be an orthogonal projection such that $BP=B$ and $PA=A$. Assume $2(k-1)\leq Tr(P)$, then,
$E_{p}\left[\|P(x-x^{*})\|^{2k}\right]$ $\leq\left(\frac{\epsilon Tr(P)}{1-\gamma}\right)^{k}.$ (29)
@@
Choose $g(x)=\|P(x-x^{*})\|^{2k}$, then if $T:=Tr(P)$ we have,
$\nabla g_{k}(x)$ $=2kg_{k-1}(x)P(x-x^{*})$ (34)
$\Delta g_{k}(x)$ $=2kTg_{k-1}(x)+4k(k-1)g_{k-1}(x)$ (35)
@@
00 0 $=E_{p}\left[\mathcal{A}g_{k}\right]$ (36)
$=E_{p}\left[\nabla g_{k}\cdot w(x)+\frac{\epsilon}{2}\Delta g_{k}\right]$ (37)
$=E_{p}\left[2kg_{k-1}P(x-x^{*})\cdot v(x)+\frac{\epsilon}{2}(2kT+4k(k-1))g_{k-1}(x)\right]$ (38)
$\leq E_{p}\left[-2k(1-\gamma)g_{k}+\frac{\epsilon}{2}(2kT+4k(k-1))g_{k-1}(x)\right]$ (39)
@@
$E_{p}\left[g_{k}\right]$ $\leq\frac{\epsilon(T+2(k-1))}{2(1-\gamma)}E_{p}\left[g_{k-1}(x)\right]$ (41)
Now apply the assumption $2(k-1)\leq T$ and the result follows by induction. ∎
@@
Assume that $\operatorname{im}A\oplus(\ker B)^{\perp}$ is only supported on the first $2r$ coordinates. The $i$th coordinate of the remainder must take the form,
@@
$|R_{i}(x)|$ $\leq M\sum_{|\alpha|=2,\alpha_{>2r}=0}\frac{2}{\alpha!}(x-x^{*})^{\alpha}$ (46)
$\leq 4rM\|P(x-x^{*})\|^{2}$ (47)
@@
$\|R(x)\|^{2}\leq 8r^{2}M\|P(x-x^{*})\|^{4}.$ (48)
Now, drop the assumption that $\operatorname{im}A\oplus(\ker B)^{\perp}$ is restricted to the first $2r$ elements.
Choosing an appropriate orthogonal matrix $Q$, it’s easy to see that $\tilde{v}(x):=Q^{T}v(Qx)$ satisfies the above assumption, with a fixed point $y^{*}=Q^{T}x^{*}$, and that if we let $y=Qx$, then
$\|v(x)-Jv(x^{*})(x-x^{*})\|=\|\tilde{v}(y)-J\tilde{v}(\tilde{y}^{*})(y-\tilde{y}^{*})\|$ (49)
so the bound follows for $v$ as well.
@@
$\leq E\left[\left\|x-x^{*}\right\|^{2}\right]^{1/2}$ (51)
@@
Fokker-Planck yields a formula for second-order moments of linear SDE stationary distribution (see for example Chapter 5.5 in [36]), such that,
00 0 $=E_{p}[(x-m)v(x)^{T}+v(x)(x-m)^{T}+\epsilon I]$ (53)
@@
$E_{p}\left[\|x-m\|\cdot\|R(x)\|\right]$ $\lesssim r\sqrt{M}E_{p}\left[(\|x-m\|\cdot\|P(x-x^{*})\|^{2}\right]$ (59)
$\lesssim r\sqrt{M}E_{p}\left[\|x-m\|^{2}\right]^{1/2}E_{p}\left[\|P(x-x^{*})\|^{4}\right]^{1/2}$ (60)
$\lesssim\frac{\epsilon r^{2}\sqrt{M}}{1-\gamma}E_{p}\left[\|x-m\|^{2}\right]^{1/2}$ (61)
$\lesssim\frac{\epsilon r^{2}\sqrt{M}}{1-\gamma}E_{p}\left[\|x-x^{*}\|^{2}+\|x^{*}-m\|^{2}\right]^{1/2}$ (62)
$\lesssim\frac{\epsilon^{3/2}r^{2}n^{1/2}\sqrt{M}}{(1-\gamma)^{2}}$ (63)
@@
$\lesssim\frac{\epsilon^{1/2}r^{2}n^{1/2}\sqrt{M}}{(1-\gamma)^{2}}$ (65)
Now, if we choose $\epsilon$ small enough to guarantee the above bound is strictly less than 1, then the Lyapunov equation given in Equation (58) has a unique solution. It’s easy to then check from the integral form of the solution that,
[…]
By assumption, $L=A\sigma^{\prime}B-I$ is Hurwitz, and in fact $\|L\|<1-\gamma$. By the fact that $\|e^{L}\|\leq e^{\|L\|}$,
[…]
$\lesssim\frac{\epsilon^{1/2}r^{2}n^{1/2}\sqrt{M}}{(1-\gamma)^{3}}$ (69)
@@
We first need a result about the equilibrium points of the drift, ruling out any degeneracies. As before, we let $x_{i}^{*}$ denote the unique zero of $v(x)+c_{i}=A\sigma(Bx)-x+c_{i}$.
@@
$J_{y}g(y,z)$ $=\sigma^{\prime}(BAy+z)BA$ (70)
$J_{z}g(y,z)$ $=\sigma^{\prime}(BAy+z)$ (71)
@@
Now, consider the random variable $y^{*}(Bc)$ where $c$ is a sampled intervention vector. Conditioning on $B$ being full-rank, $Bc$ has a density in $\mathbb{R}^{r}$, and therefore so does $y^{*}(Bc)$.

### v3

@@ 5 Experiments
To validate the theory given above, we consider experiments demonstrating the recovery of SDE parameters and generalization to unseen interventions. To show the broad applicability of the theory, we consider multiple loss functions: a loss acting directly on the parameters in the linear case (Rohbeck et al., 2024), a kernelized Stein discrepancy (Barp et al., 2019), and rollout loss for neural SDEs (Kidger et al., 2021).
@@ 5.1 Linear SDE Recovery
Because we are fitting linear SDEs where the stationary distribution is Gaussian, we can choose a very simple loss that matches the population mean and covariance of each intervented distribution, to verify the claim of identifiability (Theorem 4.4). This loss is akin to the one used in the Bicycle method introduced in Rohbeck et al. (2024) that fits interventional linear SDEs. Assuming the noise scale $\epsilon$ is known and $L$ denotes the true drift, we fit the parameters $\hat{A}$, $\hat{B}$, and $\hat{D}$. Letting our estimate for the drift be denoted by $\hat{L}:=\hat{A}\hat{B}-\hat{D}$, the loss function matches the mean and covariance of each intervention:
$\mathcal{L}_{lin}(\hat{A},\hat{B},\hat{D})=\|\hat{L}^{-1}C-L^{-1}C\|_{F}+\|\hat{L}\omega+\omega\hat{L}^{T}+\epsilon I\|_{F}.$ (8)
@@
We confirm our theory: when the decay is known, r2r-2 interventions fails to grant identifiability with poor performance and extremely high variance, even with many independent initializations, while $r$ interventions clearly succeed. When the decay isn’t known, we use a modest oversampling of $r\log(n)$ interventions. This amount is informed by the literature on matrix completion, where $\Omega(rn\log(n))$ revealed matrix entries were proven to be necessary for a rank $r$ matrix recovery problem (Candès & Tao, 2010). We scale down by a factor of $n$ because in our setting each intervention yields the mean vector of the perturbed linear SDE in $\mathbb{R}^{n}$, hence $n$ measurements.
(a) $n=100$ with fixed decay
(b) $n=200$ with fixed decay
(c) $n=100$ with learned decay
(d) $n=200$ with learned decay
Figure 2: Normalized Frobenius error of learned drift against true drift in linear SDEs with $k$ independent Gaussian interventions. Error bars are standard deviation over 5 independent runs.
5.2 Nonlinear SDE Recovery
We repeat this verification of identifiability in the setting of Theorem 4.8, using a very small value of $\epsilon=10^{-5}$ to approximately linearize. For some randomly chosen $A$ and $B$, we define the true drift $v(x)=A\sigma(Bx)-x$ and use each intervened drift $v+c_{i}$ to calculate the corresponding sample mean $m_{\epsilon}^{i}$ and sample covariance $\Sigma_{\epsilon}^{i}$.
We define a parameterized drift of the form $\hat{v}_{\theta}(x)=\hat{A}\sigma_{\theta}(\hat{B}x)-x$ where $\sigma_{\theta}$ is a learned MLP that acts elementwise. We train with the loss function
$\mathcal{L}_{nonlin}(\hat{A},\hat{B},\theta)=\sqrt{\sum_{i=1}^{k}\left\|\hat{v}_{\theta}(m_{\epsilon}^{i})+c_{i}\right\|^{2}}+\sqrt{\sum_{i=1}^{k}\left\|J\hat{v}_{\theta}(m_{\epsilon}^{i})(\Sigma_{\epsilon}^{i}/\epsilon)+(\Sigma_{\epsilon}^{i}/\epsilon)J\hat{v}_{\theta}(m_{\epsilon}^{i})^{T}+I\right\|^{2}}$ (9)
The motivation for this loss comes from Theorem 4.7. The first term enforces the approximate fixed point property of the mean, and the second term enforces the approximate Lyapunov equation of the covariance. Thus we can train $\hat{v}_{\theta}$ to match the first and second moments of the true stationary distribution of $v$, which will be approximately Gaussian for sufficiently small $\epsilon$.
To empirically evaluate identifiability, we calculate the normalized error in recovering $A$ and $B$ from the parameters $\hat{A}$ and $\hat{B}$, chosen from the run with minimum training loss across 100 independent runs. Further details are given in the Appendix. We observe that $r+1$ interventions enable consistently good recovery of the underlying drift parameters, whereas $r$ interventions leads to high variance estimates. This suggests that the $r+1$ upper bound proven in Theorem 4.8 may be tight, and offers a direction for further work refining the theory of identifying the nonlinear case.
Table 1: Normalized Frobenius error of drift parameters with $k$ independent Gaussian interventions, over 20 independent runs.
AA Error $B$ Error
$k=r$ $0.13\pm 0.13$ $0.14\pm 0.15$
$k=r+1$ $\bm{0.065\pm 0.04}$ $\bm{0.07\pm 0.04}$
@@ 5.3 Synthetic Nonlinear SDE Generalization
For an evaluation of Theorem 4.8, we apply the insight that identifiability doesn’t require knowing $\sigma$ and consider learnable MLPs. Learnable activations (Goyal et al., 2019) have been proposed before (Apicella et al., 2019) but to our knowledge not previously applied to SDE parameterizations.
To assess generalizability, we consider a loss function proposed by Lorch et al. (2024): the kernel deviation from stationarity (KDS), or equivalently the kernelized Stein discrepancy of the stationary distribution (Barp et al., 2019). When the parametric drift is defined as $v_{\theta}(x)=A\sigma(Bx)-Dx$ and the target stationary distribution is $\mu$, the KDS is,
@@
where $k$ is a given kernel and $\mathcal{A}_{x}$ is the SDE generator acting on the variable $x$ (for more details see Lorch et al. (2024)). We use samples from the true stationary distribution $\mu$ to approximate these expectations, further training details are given in the Appendix. This loss is nevertheless unstable when using noise small enough to reach the limit proven in Theorem 4.8, and therefore we assess generalizability according to performance on unseen interventions.
We observed numerical instability in the Sinkhorn divergence without large entropic regularization (Cuturi, 2013). Therefore, we follow Zhang et al. (2023); Lorch et al. (2024) and evaluate with the mean squared error (MSE) between the true and predicted distribution. Further details are given in Section B.3. The results are given in Table 2. We observe a clear benefit for low noise SDEs, which gets smaller as the noise gets bigger and hence further away from our proven settings. This is consistent with our theory, and also intuitive: as the noise gets larger, the stationary distribution gets smoother and has less dependence on the intervention vectors. Nevertheless, in the low noise regime we see that learnable activations improve prediction on test interventions without overfitting.
Table 2: Mean distance between true and predicted distribution with $n=20$, $r=3$, over 20 independent seeds with 20 test interventions per seed, with sigmoid activation versus learned activation.
@@ 5.4 Simulated GRN SDE Generalization
We consider an applied setting for learnable activations, and how they impact the recovery of GRNs. In the genomics context, $n$ is the number of expressed genes in the thousands, and $r$ corresponds to a much smaller number of latent gene modules/pathways. Because there are few fully known regulatory networks we rely on semi-synthetic data to produce data with similar characteristics to true perturbed transcriptomic samples and a ground-truth GRN.
We use the PerturbODE model from Lin et al. (2025), similar to our setup and modified to handle a loss for parameterized SDEs (Kidger et al., 2021). We consider the inferred GRN as the matrix multiplication of the weight matrices in the MLP that parameterizes the drift, corresponding to a first-order Taylor approximation of the dynamics. GRN recovery can be measured as a classification task of individual edges. As a baseline, we also consider the Bicycle interventional linear SDE method (Rohbeck et al., 2024). To simulate a known GRN, we use the boolean ODE based-simulator BEELINE (Pratapa et al., 2020). BEELINE uses the provided GRN to parametrize an SDE in the space of mRNA expression and protein expression, with physically plausible regulation dynamics. More precise details are given in the Appendix in Section B.4.
@@
Figure 3: Gene regulatory network recovery on three tested SDE models for 5 independent runs.
@@ 6 Discussion
We confirm identifiability directly in the linear case and non-linear, low-noise limit, and see improved generalization using learned activations for small noise. It may be beneficial to increase the capacity of the SDE beyond two-layer neural networks, if the extracted drift were still identifiable.
We note some limitations of the given theory. The results demonstrate identifiability in a setup with infinitesimal small noise and restrictions on the possible drift functions. Generalizing to larger noise regimes may be possible with higher order moment information about the stationary distribution, although understanding the stationary distribution outside of linearization is very challenging. Broader parameterizations of the drift or interventions are interesting extensions for future work.
@@ Appendix A Proofs of Results
and set $D=I$ and $\epsilon=1$. If we also set
@@
then it is straightforward to confirm that the drift matrix L=ABIL=AB-I satisfies the Lyapunov equation
$L\omega+\omega L^{T}+I=0$ (20)
Now, consider any skew-symmetric matrix $Q$ that is only supported in the top left $r\times r$ block. Then $AB+Q\omega^{-1}$ will only be supported in the top left block and therefore still have rank $r$. Furthermore, $Q$ obeys the trivial Lyapunov equation $Q\omega+\omega Q^{T}=0$, so we have that the drift matrix $\hat{L}:=AB+Q\omega^{-1}-I$ also satisfies the Lyapunov equation $\hat{L}\omega+\omega\hat{L}^{T}+I=0$. Choosing the norm of $Q$ small enough guarantees that $\hat{L}$ is still Hurwitz, while choosing $Q\neq 0$ guarantees $L\neq\hat{L}$.
In the worst case, every intervention is of the form $e_{i}$ for $i>r$. The block structure and Woodbury matrix identity imply that $(AB+Q\omega^{-1}-I)^{-1}e_{i}=-e_{i}$ for any $Q$ chosen as above, and therefore $\hat{L}^{-1}e_{i}=-L^{-1}e_{i}$.
We conclude that for each intervention $e_{i}$ with $i>r$, by Theorem 2.1 the drift matrices $L$ and $\hat{L}$ both induce identical stationary distributions. Thus, even with nrn-r interventions, the drift is not identifiable.
@@
Suppose the true $A$ and $B$ are sampled according to Assumption 4.2, with the constraint that a.s. $\|A\|,\|B\|\leq\sqrt{\gamma}$. Then L=ABDL=AB-D is a.s. identifiable.
@@
$0$ $=L\omega P+\omega L^{T}P+\epsilon P$ (22)
@@
$(\omega D+\epsilon I)P$ is rank nrn-r, so a.s. $\operatorname{im}(\omega D+\epsilon I)P\oplus\operatorname{im}C$ is $n$ dimensional. Hence we recover $L^{-1}$ and therefore $L$.
@@
Sample $A\in\mathbb{R}^{n\times r}$ and $B\in\mathbb{R}^{r\times n}$ according to Assumption 4.2. Then a.s. $AB$ is not identifiable with r2r-2 interventions.
@@
Assume the decay is fixed at $D=I$, so L=ABIL=AB-I with induced covariance $\omega$. With r2r-2 interventions, $ABL^{-1}C$ is at most rank r2r-2, and by assumption the kernel of $A^{T}$ is at most dimension nrn-r, so there is guaranteed to be a two dimensional subspace orthogonal to the previous spaces, say with basis vectors $u$ and $v$. Let $Q^{*}=uv^{T}-vu^{T}$.
@@
Suppose we have nr1n-r-1 interventions. Since $A$ has $r$ columns, there must be a vector $u$ in $\mathbb{R}^{n}$ orthogonal to the columns of $A$ and $C$.
@@
Consider a vector field $v(x)=A\sigma(Bx)-x+c$ with $\|A\|,\|B\|\leq 1$, $\|\sigma\|_{\infty}=\gamma<1$. Let $p$ be the stationary distribution of our usual SDE, $x^{*}$ be the unique stationary point of $v$, and $P$ be an orthogonal projection such that $BP=B$ and $PA=A$. Assume $2(j-1)\leq Tr(P)$, then,
$E_{p}\left[\|P(x-x^{*})\|^{2j}\right]$ $\leq\left(\frac{\epsilon Tr(P)}{1-\gamma}\right)^{j}.$ (30)
@@
Choose $g_{j}(x)=\|P(x-x^{*})\|^{2j}$, then if $T:=Tr(P)$ we have,
$\nabla g_{j}(x)$ $=2jg_{j-1}(x)P(x-x^{*})$ (35)
$\Delta g_{j}(x)$ $=2jTg_{j-1}(x)+4j(j-1)g_{j-1}(x)$ (36)
@@
$0$ $=E_{p}\left[\mathcal{A}g_{j}\right]$ (37)
$=E_{p}\left[\nabla g_{j}\cdot v(x)+\frac{\epsilon}{2}\Delta g_{j}\right]$ (38)
$=E_{p}\left[2jg_{j-1}P(x-x^{*})\cdot v(x)+\frac{\epsilon}{2}(2jT+4j(j-1))g_{j-1}(x)\right]$ (39)
$\leq E_{p}\left[-2j(1-\gamma)g_{j}+\frac{\epsilon}{2}(2jT+4j(j-1))g_{j-1}(x)\right]$ (40)
@@
$E_{p}\left[g_{j}\right]$ $\leq\frac{\epsilon(T+2(j-1))}{2(1-\gamma)}E_{p}\left[g_{j-1}(x)\right]$ (42)
Now apply the assumption $2(j-1)\leq T$ and the result follows by induction. ∎
@@
Assume that $\operatorname{im}A\oplus(\ker B)^{\perp}$ is only supported on the first $2r$ coordinates. The $i$th coordinate of the Taylor remainder must take the form,
@@
$|R_{i}(x)|$ $\leq M\sum_{|\alpha|=2,\alpha_{>2r}=0}\frac{2}{\alpha!}\left|(x-x^{*})^{\alpha}\right|$ (47)
$\leq M\left(\sum_{j=1}^{2r}|x_{j}-x_{j}^{*}|\right)^{2}$ (48)
$\leq 2rM\left(\sum_{j=1}^{2r}|x_{j}-x_{j}^{*}|^{2}\right)$ (49)
$=2rM\|P(x-x^{*})\|^{2}$ (50)
@@
$\|R(x)\|^{2}\leq 8r^{3}M^{2}\|P(x-x^{*})\|^{4}.$ (51)
Now, we drop the assumption that $\operatorname{im}A\oplus(\ker B)^{\perp}$ is restricted to the first $2r$ elements and extend the result more generally.
Consider the orthogonal matrix $Q$ such that $\operatorname{im}QA\oplus(\ker BQ^{T})^{\perp}$ is restricted to the first $2r$ components, then by design the above result can be applied to the new drift function
$\tilde{v}(y):=QA\sigma(BQ^{T}y)-y+Qc$ (52)
whose unique stationary point is $y^{*}=Qx^{*}$.
Hence, applying the above result to the remainder term $\tilde{R}(y):=\tilde{v}(y)-J\tilde{v}(y^{*})(y-y^{*})$, with $\tilde{P}$ the projection onto the first $2r$ components, we get
$\|\tilde{R}(y)\|\leq 2\sqrt{2}r^{3/2}M\|\tilde{P}(y-y^{*})\|^{2}$ (53)
If we let $y=Qx$, then some algebra reveals
$\tilde{R}(y)$ $=\tilde{v}(y)-J\tilde{v}(y^{*})(y-y^{*})$ (54)
$=QA\sigma(BQ^{T}y)-y+Qc-(QAJ\sigma(BQ^{T}y^{*})BQ^{T}-I)(y-y^{*})$ (55)
$=Q(A\sigma(Bx)-x+c-(AJ\sigma(Bx^{*})B-I)(x-x^{*}))$ (56)
$=Q(v(x)-Jv(x^{*})(x-x^{*}))$ (57)
$=QR(x)$ (58)
And so we can finally conclude
$\|R(x)\|$ $=\|\tilde{R}(y)\|$ (59)
$\leq 2\sqrt{2}r^{3/2}M\|\tilde{P}(y-y^{*})\|^{2}$ (60)
$=2\sqrt{2}r^{3/2}M\|Q^{T}\tilde{P}Q(x-x^{*})\|^{2}$ (61)
Note that by our choice of $Q$, $Q^{T}\tilde{P}Q$ is the orthogonal projection onto $\operatorname{im}A\oplus(\ker B)^{\perp}$, and so we have the bound.
@@
$\leq E_{p}\left[\left\|x-x^{*}\right\|^{2}\right]^{1/2}$ (63)
@@
Fokker-Planck yields a formula for second-order moments of the SDE stationary distribution (see for example Chapter 5.5 in Särkkä & Solin (2019)), such that,
$0$ $=E_{p}[(x-m)v(x)^{T}+v(x)(x-m)^{T}+\epsilon I]$ (65)
@@
$E_{p}\left[\|x-m\|\cdot\|R(x)\|\right]$ $\lesssim r^{3/2}ME_{p}\left[(\|x-m\|\cdot\|P(x-x^{*})\|^{2}\right]$ (71)
$\lesssim r^{3/2}ME_{p}\left[\|x-m\|^{2}\right]^{1/2}E_{p}\left[\|P(x-x^{*})\|^{4}\right]^{1/2}$ (72)
$\lesssim\frac{\epsilon r^{5/2}M}{1-\gamma}E_{p}\left[\|x-m\|^{2}\right]^{1/2}$ (73)
$\lesssim\frac{\epsilon r^{5/2}M}{1-\gamma}E_{p}\left[\|x-x^{*}\|^{2}+\|x^{*}-m\|^{2}\right]^{1/2}$ (74)
$\lesssim\frac{\epsilon^{3/2}r^{5/2}n^{1/2}M}{(1-\gamma)^{2}}$ (75)
@@
$\lesssim\frac{\epsilon^{1/2}r^{5/2}n^{1/2}M}{(1-\gamma)^{2}}$ (77)
Now, if we choose $\epsilon$ small enough to guarantee the above bound is strictly less than 1, then the Lyapunov equation given in Equation (70) has a unique solution. Furthermore, the Lyapunov equation has an integral form (Särkkä & Solin, 2019), which states that the unique positive definite matrix $\omega$ that satisfies $L\omega+\omega L^{T}+QQ^{T}=0$ can be written as
$\omega=\int_{0}^{\infty}e^{Lt}QQ^{T}e^{L^{T}t}dt$ (78)
It’s easy to then conclude from the integral form of the Lyapunov equation in Equation (70) that,
$\|\Sigma/\epsilon-\omega\|\lesssim\frac{\epsilon^{1/2}r^{5/2}n^{1/2}M}{(1-\gamma)^{2}}\int_{0}^{\infty}\|e^{Lt}\|\|e^{L^{T}t}\|dt.$ (79)
By assumption, $L$ is Hurwitz, and in fact $\|L\|<1-\gamma$. By the fact that $\|e^{L}\|\leq e^{\|L\|}$,
$\|\Sigma/\epsilon-\omega\|$ $\lesssim\frac{\epsilon^{1/2}r^{5/2}n^{1/2}M}{(1-\gamma)^{2}}\int_{0}^{\infty}e^{2\|L\|t}dt$ (80)
$\lesssim\frac{\epsilon^{1/2}r^{5/2}n^{1/2}M}{(1-\gamma)^{2}}\int_{0}^{\infty}e^{2(1-\gamma)t}dt$ (81)
$\lesssim\frac{\epsilon^{1/2}r^{5/2}n^{1/2}M}{(1-\gamma)^{3}}$ (82)
@@
We begin with a series of lemmas. We first need a result about the equilibrium points of the drift, ruling out any degeneracies. As before, we let $x_{i}^{*}$ denote the unique zero of $v(x)+c_{i}=A\sigma(Bx)-x+c_{i}$.
@@
$J_{y}g(y,z)$ $=J\sigma(BAy+z)BA$ (83)
$J_{z}g(y,z)$ $=J\sigma(BAy+z)$ (84)
@@
Now, condition on $A$ and $B$, which are a.s. full rank. Consider the random variable $y^{*}(Bc)$ where $c$ is a sampled intervention vector. Because $Bc$ has a density in $\mathbb{R}^{r}$, so does $y^{*}(Bc)$.